OpenStudy (anonymous):

HEELPP!! Marianna deposited $200 into her bank account at the end of each month for 8 months. The account pays 2.9% per annum, compounded annually.

OpenStudy (tkhunny):

Not quite. That is WAY too much interest. Annual Compounding of Monthly Payments requires a different sort of attention. i = 0.029 -- Annual Interest Rate r = 1 + i = 1.029 -- Annual Accumulation Factor s = r^(1/12) = 1.002385128 -- Monthly Accumulation Factor Time 0 mo $0.00 Time 1 mo $200.00 Time 2 mo $200.00 + $200.00s Time 3 mo $200.00 + $200.00s + $200.00s^2 Time 4 mo $200.00(1 + s + s^2 + s^3) Time 5 mo $200.00(1 + s + s^2 + s^3 + s^4) = \($200\cdot\dfrac{s-s^{5}}{1-s}\) Time 6 mo \($200\cdot\dfrac{s-s^{6}}{1-s}\) Time 7 mo \($200\cdot\dfrac{s-s^{7}}{1-s}\) Time 8 mo \($200\cdot\dfrac{s-s^{8}}{1-s}\)

OpenStudy (anonymous):

What formula did you use? :(

OpenStudy (anonymous):

The formula given to us is A = R[(1 + i)^n - 1]/i

OpenStudy (tkhunny):

I hate "the formula", preferring instead to CREATE it on the fly, as shown above. I did not use any formula. Nor shall I if I can possibly avoid it. I'm not sure that is the right formula. I had a typo. My numerators should be "1-", not "s-". It's easy to go wrong, formula or no. The difficulty with that formula is what is "i"? Your "i" is my "s-1". You can take it from there.